# Characterization of an MPPC-Based Scintillator Telescope and Measurement of Cosmic Muon Angular Distribution

### v1

@@ I Introduction
Earth is under constant bombardment by cosmic rays, a flux of highly energetic charged particles originating from outer space . This primary radiation, composed of approximately 89% protons, 9% helium nuclei, 1% electrons and 1% heavier nuclei, is thought to be accelerated to relativistic speeds by powerful galactic events such as supernovae [undef]. Upon entering the upper atmosphere, these primary particles collide with air molecules, initiating a cascade of lighter, secondary particles known as an extensive air shower. In these initial interactions, unstable mesons—primarily pions ($\pi$) and kaons ($\kappa$)—are produced. The charged pions are of particular importance, as they decay almost instantaneously into the particles that are the focus of this study: muons [undefa] refer figure 1.
@@ II-A1 Muon Interaction with Matter and Energy Loss
When this formula is analyzed as a function of the incident particle’s momentum, a distinct pattern emerges, as illustrated in Figure 3. The energy loss decreases as energy increases until it reaches a specific minimum point, after which the curve becomes relatively constant. Particles moving with momenta near this minimum are referred to as Minimum Ionizing Particles (MIPs).
@@ II-A2 Scintillation Detection with Multi-Pixel Photon Counters (MPPCs)
The emitted photons are detected by a Multi-Pixel Photon Counter (MPPC), also known as a Silicon Photomultiplier (SiPM), coupled to the scintillator. These devices consist of a dense array of thousands of microscopic avalanche photodiodes (APDs) connected in parallel. To detect the faint light signals, the MPPC is operated in “Geiger mode,” achieved by applying a reverse bias voltage slightly above the breakdown voltage. In this state, a single photon hitting a pixel triggers a massive “avalanche” of electrons.
Each pixel within the MPPC acts like a binary switch: if one photon hits, one pixel fires; if one hundred photons hit, one hundred pixels fire simultaneously. The final output current is the sum of all firing pixels, producing a measurable electrical pulse large enough for the oscilloscope to detect. However, thermal agitation can sometimes trigger an avalanche even without incident light, creating “dark noise,” but these low-amplitude signals are easily filtered out by the trigger threshold. For a detailed description of the MPPC operation principles and noise characteristics, see [undefe].
@@ II-A3 Electrical Characteristics of MPPCs
Figure 4: Photograph of the Hamamatsu MPPC sensors used in the experiment.
@@
In this equation, $V_{\text{operation}}$ is the bias voltage applied to the sensor, and $V_{\text{breakdown}}$ is the threshold voltage necessary to start the avalanche process. The difference between these terms, $(V_{\text{operation}}-V_{\text{breakdown}})$, is commonly referred to as the overvoltage. With $C_{\text{pixel}}$ representing the pixel capacitance, the equation shows that the Gain ($G=Q_{0}/e$) scales linearly with this overvoltage. Therefore, increasing the operation voltage strictly increases the charge output, providing a stronger signal for detection [undeff].
II-A3b Dark Count Rate
However, this high sensitivity comes with a side effect. Inside the MPPC, pulses are not only created by light. They can also be triggered by heat energy, which knocks electrons loose inside the silicon chip. These false signals are called ”dark pulses” because they happen even when the detector is in total darkness.
The problem is that these dark pulses often look exactly like the real signals caused by photons. We cannot tell them apart just by looking at their shape. The number of these false pulses per second is called the Dark Count Rate. Because these pulses come from thermal energy, the rate depends heavily on the temperature of the sensor. The formula describing this relationship is [undefg]:
@@ II-A4 The Coincidence Technique for Background Rejection
In the previous section, we discussed setting a trigger threshold to reduce noise. However, relying on a single detector still has limitations, as it can pick up various background signals (e.g., dark noise, natural radiation, electronic interference), creating uncertainty about whether a trigger is a genuine muon or random noise. To overcome this, we arrange multiple detectors in a vertical alignment, creating a ”muon telescope.” This setup operates on an AND logic, recording an event only when all detectors fire simultaneously, a standard technique for signal selection in nuclear physics [undefd].
@@
However, since real electronics are not infinitely fast, we must define a short time limit known as the “coincidence window” ($\tau$). Occasionally, random noise pulses might overlap within this window just by luck; these are called “accidental coincidences.” For a three-detector system, the rate of these accidents is given by [undefh] :
@@ II-B Models for Muon Angular Distribution
In this section, we examine four different models to see which one best describes the data collected in our experiment. We start with the standard $\cos^{2}\theta$ approximation, which is the most common textbook description of muon flux. We then look at the General $\cos^{n}\theta$ model used by Pethuraj et al [undefi], which treats the steepness of the curve as a variable rather than a fixed number.
We also consider a more complex physical model by Shukla and Sankrith[undefj], which accounts for the curvature of the Earth, and finally, the Schwerdt model [undefk]. The Schwerdt model is an empirical formula and includes extra terms to account for background offsets. The following subsections explain the physics and mathematics behind each of these approaches.
@@ II-B1 The $\cos^{2}(\theta)$ Approximation
II-B2 The General $\cos^{n}(\theta)$ Model (Pethuraj et al.)
@@
To account for these variations, Pethuraj et al.[undefi] proposed leaving the exponent as a free parameter ($n$) rather than fixing it at exactly 2. The model is expressed as:
@@
In the study by Pethuraj et al. [undefi], they found a value of $n\approx 2.00\pm 0.04$ at an altitude of $\sim 160$ m. This confirms that the distribution is indeed close to the theoretical prediction but allows for experimental fine-tuning.
II-B3 The Shukla & Sankrith Model
@@
Shukla and Sankrith[undefj] derived a more accurate formula that accounts for the curvature of the Earth. Instead of a simple flat layer, they model the atmosphere as a spherical shell. Their model replaces the simple cosine term with a path-length factor $D(\theta)$. The formula is:
@@ II-B4 The Schwerdt Model
with $a$ representing a vertical stretch, $b$ a horizontal stretch, $c$ a horizontal stretch, and $d$ a vertical shift.[undefk]
@@ II-C Apparatus
The complete experimental apparatus used for this study is shown in Figure 2. The system integrates the scintillation detectors, signal processing electronics, and data acquisition hardware into a compact tabletop assembly. The specific components and their configuration are detailed in the following subsections.
@@ II-C1 Detector Components: Scintillators and MPPCs
When a cosmic muon passes through one of these bars, it interacts with the plastic to produce a faint flash of blue light [undefd]. However, since our light sensor is very small compared to the size of the bar, simply attaching it to the plastic would result in a significant loss of signal. To capture this light efficiently, we embedded a wavelength-shifting fiber (see Figure 6), about 1 mm in diameter, along the center of each scintillator. This fiber performs a critical function: it absorbs the original blue light and re-emits it as green light. This green color is actually visible to the naked eye at the tips of the fiber. By shifting the color, the fiber acts as a light guide, trapping the signal inside and funneling it directly to the end of the bar where our sensor is mounted [undefd].
@@
The device used to read the optical signal is the Hamamatsu S10362-11 Series Multi-Pixel Photon Counter (MPPC) [undefl]. This device acts as the modern equivalent of a Photomultiplier Tube (PMT). While traditional PMTs are bulky, fragile, and require high voltages, this MPPC is a compact silicon device that is physically robust and operates at a relatively low bias voltage (approximately 70 V).
The specific sensor employed in our experiment has a tiny effective active area of 1 mm by 1 mm. Despite its small size, it contains a dense array of avalanche photodiodes (pixels). Depending on the specific model variation used, the sensor contains between 100 and 1600 individual pixels. These pixels operate in Geiger mode, effectively acting as binary switches that fire when struck by a photon.
According to the manufacturer specifications [undefl], the peak sensitivity of this sensor is at 440 nm. However, its spectral response range is broad (320 to 900 nm), which allows it to efficiently detect the green light emitted by the wavelength-shifting fiber [undefe]. This combination of spectral matching and high internal gain (ranging from $10^{5}$ to $10^{6}$) ensures that even the faint light pulses from a single muon are amplified enough to be recorded by our electronics.
@@
To operate the MPPCs, a custom circuit board was designed to apply the necessary bias voltage and extract the signal. As shown in Figure 7, the circuit uses a network of resistors and capacitors to connect the power supply to the sensor.
@@ II-C2 Data Acquisition System and Triggering Logic
In the past, analyzing these types of nuclear signals required racks of specialized and expensive digitizing hardware. However, digital technology has evolved to the point where a single, high-performance desktop oscilloscope can now replace those bulky devices [undeff]. Modern scopes offer advanced features that allow them to act as both the eyes of the experiment, visualizing the pulses, and the brain, performing the logic required to identify valid data.
For this specific setup, we selected the RIGOL DHO924 (see Figure 8). This instrument represents a significant improvement over standard laboratory equipment because it is a high-resolution 12-bit oscilloscope. Most common scopes only offer 8-bit resolution, which can make small signals look grainy or pixelated. The 12-bit resolution of our device provides 16 times more vertical precision, allowing us to see the signal with exceptional clarity. This precision allows us to set a distinct voltage threshold to reject low-amplitude thermal noise, ensuring we only analyze valid signals. Technical specifiations are given at table I.
@@ II-C3 The Three-Fold Coincidence Telescope Geometry
The geometry of the stack—specifically the physical size of the detectors ($25~\text{cm}\times 2.5~\text{cm}$) and the vertical distance between the top and bottom counters—determines the solid angle of acceptance. This geometry essentially restricts the detector’s field of view. A tall telescope (large separation) sees a small patch of sky, while a compact telescope (small separation) observes a wider area. Since the frame is rigid, this geometric factor remains constant for all angular measurements, ensuring that the comparison between the $0^{\circ}$ flux and the $90^{\circ}$ flux is scientifically valid.
@@ III-B Threshold Optimization and Selection
As the table shows, the singles rates are very high at the 1 and 2 p.e. levels. These high rates of random noise are the main reason we would get false “accidental” coincidences.
@@
Next, we measured the actual two-fold and three-fold coincidence rates, with the results shown in Table V. A representative waveform showing the simultaneous firing of three detector channels as observed on the oscilloscope screen is shown in Figure 10.
@@ IV-A Data Collection
To establish a high-precision baseline for the vertical muon flux, an extended 12-hour measurement was conducted with the detector positioned at a $0^{\circ}$ zenith angle. For all other angles, from $10^{\circ}$ to $90^{\circ}$ in $10^{\circ}$ increments, the number of coincidence events was recorded for a total duration of one hour each. The data for all runs was binned into 5-minute intervals to monitor the stability of the count rate. Throughout the data acquisition period, the detector was maintained in complete darkness using a light-proof black sheet (see Figure 11) to ensure signal integrity.
@@ IV-B Data Stability and Final Rate Calculation
Before the final analysis, the stability of the data for each run was evaluated. The 1-hour and 24-hour measurements were binned into 5-minute intervals, and the number of counts in each interval was plotted against time to check for drifts or anomalies. A representative stability plot for the long 24-hour measurement at 0 degrees is shown in Figure 12.
@@
With the stability of the runs verified, a single average rate and its statistical error were calculated for each angle by dividing the total number of events by the total time. The 0-degree point, benefiting from the much longer run time, has a significantly smaller statistical uncertainty. These final, processed rates are the basis for the main analysis in the following section.
@@ IV-C Angular Dependence of the Muon Rate
$\theta$ () Hits Rate (Hz) Err (Hz) Flux
0 [N] 0.152 0.0065 2.43
10 [N] 0.142 0.0063 2.27
20 [N] 0.131 0.0060 2.09
30 [N] 0.108 0.0055 1.72
40 [N] 0.088 0.0049 1.40
50 [N] 0.077 0.0046 1.24
60 [N] 0.048 0.0037 0.77
70 [N] 0.029 0.0028 0.46
80 [N] 0.017 0.0022 0.27
@@
To visualize the relationship between the muon rate and the zenith angle, this data is plotted in Figure 13.
@@
The data in Figure 13 clearly demonstrates a strong dependence on the zenith angle. The measured rate is at its maximum at $0^{\circ}$ and decreases monotonically as the angle increases towards $90^{\circ}$. This trend is qualitatively consistent with the expected angular distribution [undefb]. A detailed quantitative analysis comparing this data to several theoretical models is performed in the next section.
@@ IV-D Comparative Analysis of Angular Distribution Models
The best-fit curve for each of these four models is shown overlaid on the experimental data in Figure 14.

### v2

@@ I Introduction
Earth is under constant bombardment by cosmic rays, a flux of highly energetic charged particles originating from outer space . This primary radiation, composed of approximately 89% protons, 9% helium nuclei, 1% electrons and 1% heavier nuclei, is thought to be accelerated to relativistic speeds by powerful galactic events such as supernovae [undef]. Upon entering the upper atmosphere, these primary particles collide with air molecules, initiating a cascade of lighter, secondary particles known as an extensive air shower. In these initial interactions, unstable mesons-primarily pions ($\pi$) and kaons ($\kappa$)-are produced. The charged pions are of particular importance, as they decay almost instantaneously into the particles that are the focus of this study: muons [undefa] refer Fig. 1.
@@ II-A1 Muon Interaction with Matter and Energy Loss
When this formula is analyzed as a function of the incident particle’s momentum, a distinct pattern emerges, as illustrated in Fig. 3. The energy loss decreases as energy increases until it reaches a specific minimum point, after which the curve becomes relatively constant. Particles moving with momenta near this minimum are referred to as Minimum Ionizing Particles (MIPs).
@@ II-A2 Scintillation Detection with Multi-Pixel Photon Counters (MPPCs)
The emitted photons are detected by an MPPC, also known as a Silicon Photomultiplier (SiPM), coupled to the scintillator. These devices consist of a dense array of thousands of microscopic Avalanche Photodiodes (APDs) connected in parallel. To detect the faint light signals, the MPPC is operated in “Geiger mode”, achieved by applying a reverse bias voltage slightly above the breakdown voltage. In this state, a single photon hitting a pixel triggers a massive “avalanche” of electrons.
Each pixel within the MPPC acts like a binary switch: if one photon hits, one pixel fires; if one hundred photons hit, one hundred pixels fire simultaneously. The final output current is the sum of all firing pixels, producing a measurable electrical pulse large enough for the oscilloscope to detect. However, thermal agitation can sometimes trigger an avalanche even without incident light, creating “dark noise”, but these low-amplitude signals are easily filtered out by the trigger threshold. Detailed description of the MPPC operation principles and noise characteristics is provided in  [undefe].
@@ II-A3 Electrical Characteristics of MPPCs
Figure 4: Hamamatsu MPPC sensors used in the experiment.  [undeff]
@@
In this equation, $V_{\text{operation}}$ is the bias voltage applied to the sensor, and $V_{\text{breakdown}}$ is the threshold voltage necessary to start the avalanche process. The difference between these terms, $(V_{\text{operation}}-V_{\text{breakdown}})$, is commonly referred to as the overvoltage. With $C_{\text{pixel}}$ representing the pixel capacitance, the equation shows that the Gain ($G=Q_{0}/e$) scales linearly with this overvoltage. Therefore, increasing the operation voltage strictly increases the charge output, providing a stronger signal for detection [undefg].
II-A3b Dark Count Rate (DCR)
However, this high sensitivity comes with a side effect. Inside the MPPC, pulses are not only created by light. They can also be triggered by heat energy, which knocks electrons loose inside the silicon chip. These false signals are called “dark pulses” because they happen even when the detector is in total darkness.
The problem is that these dark pulses often look exactly like the real signals caused by photons. We cannot tell them apart just by looking at their shape. The number of these false pulses per second is called the Dark Count Rate (DCR). Because these pulses come from thermal energy, the rate depends heavily on the temperature of the sensor [undefh]. The formula describing this relationship is given below
@@ II-A4 The Coincidence Technique for Background Rejection
In the previous section, we discussed setting a trigger threshold to reduce noise. However, relying on a single detector still has limitations, as it can pick up various background signals (e.g., dark noise, natural radiation, electronic interference), creating uncertainty about whether a trigger is a genuine muon or random noise. To overcome this, we arrange multiple detectors in a vertical alignment, creating a “muon telescope”. This setup operates on an AND logic, recording an event only when all detectors fire simultaneously, a standard technique for signal selection in nuclear physics [undefd].
@@
However, since real electronics are not infinitely fast, we must define a short time limit known as the “coincidence window” ($\tau$). Occasionally, random noise pulses might overlap within this window just by luck; these are called “accidental coincidences”. For a three-detector system  [undefi], the rate of these accidents is given by
@@ II-B Models for Muon Angular Distribution
In this section, we examine four different models to see which one best describes the data collected in our experiment. We start with the standard $\cos^{2}\theta$ approximation, which is the most common textbook description of muon flux. We then look at the General $\cos^{n}\theta$ model used by Pethuraj et al [undefj], which treats the steepness of the curve as a variable rather than a fixed number.
We also consider a more complex physical model by Shukla and Sankrith[undefk], which accounts for the curvature of the Earth, and finally, the Schwerdt model [undefl]. The Schwerdt model is an empirical formula and includes extra terms to account for background offsets. The following subsections explain the physics and mathematics behind each of these approaches.
@@ II-B1 The $\cos^{2}(\theta)$ Approximation
II-B2 The General $\cos^{n}(\theta)$ Model (Pethuraj et al. [undefj])
@@
To account for these variations, Pethuraj et al. proposed leaving the exponent as a free parameter ($n$) rather than fixing it at exactly 2. The model is expressed as:
@@
In the study by Pethuraj et al. [undefj], they found a value of $n\approx 2.00\pm 0.04$ at an altitude of $\sim 160$ m. This confirms that the distribution is indeed close to the theoretical prediction but allows for experimental fine-tuning.
II-B3 The Shukla & Sankrith Model [undefk]
@@
Shukla and Sankrith derived a more accurate formula that accounts for the curvature of the Earth. Instead of a simple flat layer, they model the atmosphere as a spherical shell. Their model replaces the simple cosine term with a path-length factor $D(\theta)$. The formula is:
@@ II-B4 The Schwerdt Model [undefl]
with $a$ representing a vertical stretch, $b$ a horizontal stretch, $c$ a horizontal stretch, and $d$ a vertical shift. [undefl]
@@ II-C Apparatus
The complete experimental apparatus used for this study is shown in Fig. 2. The system integrates the scintillation detectors, signal processing electronics, and data acquisition hardware into a compact tabletop assembly. The specific components and their configuration are detailed in the following subsections.
@@ II-C1 Detector Components: Scintillators and MPPCs
When a cosmic muon passes through one of these bars, it interacts with the plastic to produce a faint flash of blue light [undefd]. However, since our light sensor is very small compared to the size of the bar, simply attaching it to the plastic would result in a significant loss of signal. To capture this light efficiently, we embedded a wavelength-shifting fiber (see Fig. 6), about 1 mm in diameter, along the center of each scintillator. This fiber performs a critical function: it absorbs the original blue light and re-emits it as green light. This green color is actually visible to the naked eye at the tips of the fiber. By shifting the color, the fiber acts as a light guide, trapping the signal inside and funneling it directly to the end of the bar where our sensor is mounted [undefd].
@@
The device used to read the optical signal is the Hamamatsu S10362-11 Series Multi-Pixel Photon Counter (MPPC) [undeff]. This device acts as the modern equivalent of a Photomultiplier Tube (PMT). While traditional PMTs are bulky, fragile, and require high voltages, this MPPC is a compact silicon device that is physically robust and operates at a relatively low bias voltage (approximately 70 V).
The specific sensor employed in our experiment has a tiny effective active area of $1~\text{mm}\times 1~\text{mm}$. Despite its small size, it contains a dense array of avalanche photodiodes (pixels). Depending on the specific model variation used, the sensor contains between 100 and 1600 individual pixels. These pixels operate in Geiger mode, effectively acting as binary switches that fire when struck by a photon.
According to the manufacturer specifications [undeff], the peak sensitivity of this sensor is at 440 nm. However, its spectral response range is broad (320 to 900 nm), which allows it to efficiently detect the green light emitted by the wavelength-shifting fiber [undefe]. This combination of spectral matching and high internal gain (ranging from $10^{5}$ to $10^{6}$) ensures that even the faint light pulses from a single muon are amplified enough to be recorded by our electronics.
@@
To operate the MPPCs, a custom circuit board was designed to apply the necessary bias voltage and extract the signal. As shown in Fig. 7, the circuit uses a network of resistors and capacitors to connect the power supply to the sensor.
@@ II-C2 Data Acquisition System and Triggering Logic
In the past, analyzing these types of nuclear signals required racks of specialized and expensive digitizing hardware. However, digital technology has evolved to the point where a single, high-performance desktop oscilloscope can now replace those bulky devices [undefg]. Modern scopes offer advanced features that allow them to act as both the eyes of the experiment, visualizing the pulses, and the brain, performing the logic required to identify valid data.
For this specific setup, we selected the RIGOL DHO924 (see Fig. 8). This instrument represents a significant improvement over standard laboratory equipment because it is a high-resolution 12-bit oscilloscope. Most common scopes only offer 8-bit resolution, which can make small signals look grainy or pixelated. The 12-bit resolution of our device provides 16 times more vertical precision, allowing us to see the signal with exceptional clarity. This precision allows us to set a distinct voltage threshold to reject low-amplitude thermal noise, ensuring we only analyze valid signals. Technical specifiations are given at Table I.
@@ II-C3 The Three-Fold Coincidence Telescope Geometry
The geometry of the stack-specifically the physical size of the detectors ($25~\text{cm}\times 2.5~\text{cm}$) and the vertical distance between the top and bottom counters-determines the solid angle of acceptance. This geometry essentially restricts the detector’s field of view. A tall telescope (large separation) sees a small patch of sky, while a compact telescope (small separation) observes a wider area. Since the frame is rigid, this geometric factor remains constant for all angular measurements, ensuring that the comparison between the $0^{\circ}$ flux and the $90^{\circ}$ flux is scientifically valid.
@@ III-B Threshold Optimization and Selection
As the Table IV shows, the singles rates are very high at the 1 and 2 p.e. levels. These high rates of random noise are the main reason we would get false “accidental” coincidences.
@@
Next, we measured the actual two-fold and three-fold coincidence rates, with the results shown in Table V. A representative waveform showing the simultaneous firing of three detector channels as observed on the oscilloscope screen is shown in Fig. 10.
@@ IV-A Data Collection
To establish a high-precision baseline for the vertical muon flux, an extended 12-hour measurement was conducted with the detector positioned at a $0^{\circ}$ zenith angle. For all other angles, from $10^{\circ}$ to $90^{\circ}$ in $10^{\circ}$ increments, the number of coincidence events was recorded for a total duration of one hour each. The data for all runs was binned into 5-minute intervals to monitor the stability of the count rate. Throughout the data acquisition period, the detector was maintained in complete darkness using a light-proof black sheet (see Fig. 11) to ensure signal integrity.
@@ IV-B Data Stability and Final Rate Calculation
Before the final analysis, the stability of the data for each run was evaluated. The 1-hour and 24-hour measurements were binned into 5-minute intervals, and the number of counts in each interval was plotted against time to check for drifts or anomalies. A representative stability plot for the long 24-hour measurement at $0^{\circ}$ is shown in Fig. 12.
@@
With the stability of the runs verified, a single average rate and its statistical error were calculated for each angle by dividing the total number of events by the total time. The $0^{\circ}$ point, benefiting from the much longer run time, has a significantly smaller statistical uncertainty. These final, processed rates are the basis for the main analysis in the following section.
@@ IV-C Angular Dependence of the Muon Rate
$\theta$ () Rate (Hz) Err (Hz) Flux
0 0.152 0.0065 2.43
10 0.142 0.0063 2.27
20 0.131 0.0060 2.09
30 0.108 0.0055 1.72
40 0.088 0.0049 1.40
50 0.077 0.0046 1.24
60 0.048 0.0037 0.77
70 0.029 0.0028 0.46
80 0.017 0.0022 0.27
@@
To visualize the relationship between the muon rate and the zenith angle, this data is plotted in Fig. 13.
@@
The data in Fig. 13 clearly demonstrates a strong dependence on the zenith angle. The measured rate is at its maximum at $0^{\circ}$ and decreases monotonically as the angle increases towards $90^{\circ}$. This trend is qualitatively consistent with the expected angular distribution [undefb]. A detailed quantitative analysis comparing this data to several theoretical models is performed in the next section.
@@ IV-D Comparative Analysis of Angular Distribution Models
The best-fit curve for each of these four models is shown overlaid on the experimental data in Fig. 14.